OpenStudy (anonymous):

What is the solution set for |2x – 3| = 13?

OpenStudy (karatechopper):

u will set up two cases for this

OpenStudy (karatechopper):

-13=2x-3 2x-3=13

OpenStudy (anonymous):

-8 AND 8 Add 3 to get 16 divide by two to get 8. Since it's absolute value it's also negative 8

OpenStudy (anonymous):

2x-3=13 and -(2x-13)=13

OpenStudy (anonymous):

|2x – 3| = 13 |2x – 3| can be 2x-3 or -2x + 3 a) 2x-3 = 13 => 2x = 16 => x = 8 b) -2x + 3 = 13 => -2x = 10 => x = -5